OpenStudy (anonymous):

Solve the system by substitution method 1/2x+1/5y=-9/5 and 1/2x+3y=1

OpenStudy (anonymous):

The answer choices i have are... A. The solution set is ( , ) type an ordered pair. B. There are infinetly many solutions C.The solution is an empty set.

myininaya (myininaya):

do you want to get rid of those yucky fractions? multiply one equation by 10 and the other by -10 then add the equations together

OpenStudy (anonymous):

i got the solution is an empty set is that correct? or did i go wrong somewhere?

myininaya (myininaya):

no i got a solution

OpenStudy (anonymous):

OH what did you get... i must have gone wrong somewhere

myininaya (myininaya):

multiply first equation by 10 multiply second equation by -10 then add the equations together

myininaya (myininaya):

10(1/2x+1/5y=-9/5) 5x+2y=-18 -10(1/2x+3y=1) -5x-30y=-10 ______________ 0x-28y=-28 -28y=-28 hope this helps

myininaya (myininaya):

so what does y=?

OpenStudy (anonymous):

y=-28 right? i think it is infinite solutions

myininaya (myininaya):

no we have -28y=-28 so y has to be 1

myininaya (myininaya):

-28(1)=-28 right?

myininaya (myininaya):

if y=1 and 5x+2y=-18, then 5x+2(1)=-18 <- solve this equation for x

OpenStudy (anonymous):

I thought a negative times a postive was a postive?

myininaya (myininaya):

negative times a positive is certainly a negative

OpenStudy (anonymous):

oh my bad i always get them confused

OpenStudy (anonymous):

is the answer (-4,-1)?

OpenStudy (anonymous):

is the answer (-4,-1)?